INAF queue observing, 2016 Jan 29-30

Observer: F. Cusano, A. Rossi
Telescope Operator: S. Allanson
Support Astronomer: Olga Kuhn
Instrument: LBC

Summary

At the sunset clouds cover about 60% of the sky . At 01:05 the sky is almost clear.

Night duration (12\xB0 el twilight) 19:20-6:10 ( 02:20-13:10 UT) = 10 h 50m

Observing time: * 10h* 20m (96%)

Weather time loss: 0.5 h (4%)

technical time loss * h* m (%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Darks     5nexpX4ditX3ndit LUCI2 for Antoniucci telluric
Dotto ID433939 LBC-b(V,B,g) LBC-r(R,z,r,I) 40m passing clouds at the beginning
Raimondo UmaII LBC-b(SDT_U,B) LBC-r(r) 1.5h  
Marconi IZW18 LBC-b(V), LBC-r(I) 10m  
Guaita M1149 Br_gamma, H2 0.5, 0.5  
DeBarros M103GLASS2c zJ 2.5h passing cirrus, 10min not good

Details

01:48 starting with LBCs the program of Dotto (some clouds at horizon)

01:52 dofpia, some problem with dofpia we are giving correction by hand

02:15 the sky is clear now, running copointing

02:23 starting science with object ID190166, we launch the script 190166_long.ob
  • dimm seeing 0.9"
  • some clouds passing through at 2:37
  • clouds more dense at 02:45 - 03:00 just half of images are probably useful
  • DIMM cannot find bright star

02:55 move to object ID 433939, script 433939_long.ob
  • still pasisng cluds , try copointing and dofpia anyway
  • 03:05 start science OB, still some cirrus, DIMM still not working
  • 3:33 dimm seeing 1", the sky looks better now

03:44 slew to copointing of ID190166
  • dofpia

03:53 send science ob 190166_long.ob of ID190166
  • dimm seeing 1.1", sky clear
  • 4:03 seeing 0.7 on R-band image
  • 4:15 dimm seeing 0.9" no clouds

04:29 slew to copointing of ID 433939
  • dofpia

4:35 send science ob 433939_long.ob of ID 433939
  • seeing 1" at 4:50 on LBC-red r-band image
  • LBC-red image 045837 might be not useful, please check. It is an unknown problem.

05:09 we slew to Raimondo, UmaII
  • run dofpia on copointing UMaII _f1_UBr_point.ob
  • dofpia, lbcrangebal
  • 05:19 start OB UMaII _f1_UBr
  • 5:30 seeing 1" on LBC-red
  • 6:00 seeing 1" on LBC-red
  • 6:10 run dofpia on UMaII _f2_UBr.ob
  • 6:12 run UMaII _f2_UBr.ob
  • 6:30 seeing 1" on LBC-red
  • 6:40 seeing 0.7" on LBC-blue

07:02 we slew to Marconi, IZW18
  • dofpia
  • 07:08 run ob science IZw18_5
  • 7:15 seeing 0.8" on LBC-red

07:20 we change to LUCI

07:46 send Guaita_BrGam
  • error committing the instrument setup
  • solved pushing initialize all_unit
  • seeing 0.8" on the guider
  • ima luci2.20160130.0007-36
  • seeing 0.8" on image at 08:14
  • seeing 0.8" on image at 08:30

08:35 send Guaita_H2
  • ima luci2.20160130.0037-66
  • seeing 0.7" on guider at 08:37
  • seeing 0.7" on image, 1.0" on guider at 09:01
  • seeing 0.8" on guider at 09:20

09:23 slew to telluric for DeBarros
  • acq luci2.20160130.0067-69 acquisition stack during check of the slit. click go failed. reset and reload + skip to item+go worked.
  • 09:32 spe luci2.20160130.70-77
  • seeing 0.9 on guider at 09:35

09:40 slew to DeBarros M103GLASS2c
  • acq luci2.20160130.0078-0081
  • spec luci2.20160130.0082-0113
  • seeing 1.2" on the guider at 10:00
  • seeing 0.8" on the guider at 10:23
  • seeing 1.1" at 10:41
  • seeing 1.2" at 10:45
  • 10:53 clouds coming in, guide star fainting, spectra 0092 affected by clouds, 0092 not good
  • we pause the script at 10:57
  • 11:14 we start the script, clous passed seeing 0.9" on the guider
  • 11:38 seeing 0.9"
  • 11:58 seeing 0.8"
  • some clouds coming in at 12:26, guide star still ok
  • 12:35 seeing 0.7", but passing cirrus, guide star ok
  • 13:00 guide star fainting, the clouds cover all the sky, 0111 not good, stop script
  • 13:10 starting again integration guide star good, seeing 0.7"
  • 13:23 end of the night

13:33 running Bias with LBCs
  • bias LBC-red 135956 with same problem that the one at LBC-red science image 045837 at ~4:35. It is a read-out problem.

13:43 running MODs calibration
  • slitflats.cal mods1b.20160130.0001-0006, mods1r.20160130.0001-0004 slit 0.6?? too tired
  • slitflats_5as. cal mods1b.20160130.0007-00, mods1r.20160130.0001-0004 5arcsec slit dual grating

During night Darks for Antoniucci telluric 4dit*3ndit LIR int
Topic revision: r24 - 01 Feb 2016, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback